Grade 11Mechanics

wave length of lazer light 6463 angastrom express in nm and microm

shubham singh
8 Years agoGrade 11

1 Answer

Samyak
8 Years ago
1 Angstrom = 10^(-10)m1nanometers=10^(-9)m1micron= 10^(-6)mNow wavelength = 6463 A°Thus in nanometers it is 646.3In micron it is .6463